OpenStudy (anonymous):

Two dice are thrown simultaneously. Given that sum of the numbers is NOT more than 5, what is the probability that sum is more than 3? 7/10 3/10 5/6 3/4 9/10

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

we are given that the sum is NOT more than 5 so this is our sample space how many ways can you get a sum NOT more than 5 ?

ganeshie8 (ganeshie8):

the sum can be 2, 3, 4 or 5, yes ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

how many ways can you get a sum of 2 ?

OpenStudy (anonymous):

one way?

ganeshie8 (ganeshie8):

yes, ``` sum = 2 (1 way) ```

ganeshie8 (ganeshie8):

what about the sum of 3, how many ways ?

OpenStudy (anonymous):

one way??

ganeshie8 (ganeshie8):

you can have : 1 + 2 or 2 + 1 so two ways right

OpenStudy (anonymous):

ohh right, i didn't think you could count both

OpenStudy (anonymous):

ways

ganeshie8 (ganeshie8):

ok good, ``` sum = 2 (1 way) sum = 3 (2 way) ```

ganeshie8 (ganeshie8):

what about sum = 4, how many ways ?

OpenStudy (anonymous):

3? 2+2, 3+1, 1+3

ganeshie8 (ganeshie8):

Very good : 1 + 3 = 4 2 + 2 = 4 3 + 1 = 4

ganeshie8 (ganeshie8):

so we have ``` sum = 2 (1 way) sum = 3 (2 way) sum = 4 (3 way) ```

ganeshie8 (ganeshie8):

can you guess how many ways there will be for getting a sum of 5 ?

OpenStudy (anonymous):

4?

ganeshie8 (ganeshie8):

thats it, we're almost done ``` sum = 2 (1 way) sum = 3 (2 way) sum = 4 (3 way) sum = 5 (4 way) ```

ganeshie8 (ganeshie8):

so how many total ways are there for getting a sum NOT more than 5 ?

OpenStudy (anonymous):

3 ways?

OpenStudy (anonymous):

oh wait...4?

ganeshie8 (ganeshie8):

nope, just add them up : 1+2+3+4 = ?

OpenStudy (anonymous):

10!

ganeshie8 (ganeshie8):

Yes! how many of them have sum more than 3 ?

OpenStudy (anonymous):

6?

ganeshie8 (ganeshie8):

how ?

ganeshie8 (ganeshie8):

just look at the table that we have created earlier ``` sum = 2 (1 way) sum = 3 (2 way) sum = 4 (3 way) sum = 5 (4 way) ```

ganeshie8 (ganeshie8):

simply add the number of ways that we can get sum of 4 and 5

OpenStudy (anonymous):

2?

ganeshie8 (ganeshie8):

how ?

OpenStudy (anonymous):

im sorry! i suck

ganeshie8 (ganeshie8):

sum = 2 (1 way) sum = 3 (2 way) `sum = 4 (3 way)` `sum = 5 (4 way)` add them : 3+4 = 7 ways

ganeshie8 (ganeshie8):

outcomes in favor = 7 total outcomes = 10 probability = ?

OpenStudy (anonymous):

right...im so stupid 7/10

ganeshie8 (ganeshie8):

you're not stupid, you're doing great! thest things takes practice

OpenStudy (anonymous):

thank you! no wonder you have so many fans! haha

ganeshie8 (ganeshie8):

haha that is because i have been here so long yw :)